Let G(x) = x5

+ x4

+ x + 1 be the polynomial generator(pattern) and Message(M) = 

11100011. Find the CRC for this message and the codeword.


1
Expert's answer
2021-09-27T10:16:56-0400

Given polynomial is

"G(x) = x^5 +x^4+x+1"

Message from the destination node"(T)=11100011"

The Divisor polynomial bit "=1\times x^5 +1\times x^4+0\times x^3+0\times x^2+1\times x+1\times1"

"=110011"

Bits to be appended to the message = (Divisor polynomial bits -1)"=6-1= 5"

Now, append 5 zero to the message bit, hence"=1110001100000"

"1110001100000\/110011" Remainder =1000

Hence, CRC = 1110

Hence, we have to append 1110 to the frame.

So, the receiver will receive 111000111110